# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ 4.2 Choosing the Uncertainty Radius
Firstly, suppose we are considering Huber’s contamination model $\mathcal{P}(P;\epsilon_{0})$ for some $\epsilon_{0}\in[0,1]$. Then $\theta$ should be chosen such that the KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ contains $\mathcal{P}(P;\epsilon_{0})$. The following result, proved in Section A.6, shows how to achieve this given an upper bound on the Stein kernel $\tau_{\infty}=\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})$.
@@
Suppose $k$ satisfies the conditions in Lemma 2 and $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$. Let $\epsilon_{0}\in[0,1]$. Then $\mathcal{P}(P;\epsilon_{0})\subseteq\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ if $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ and this bound is tight, i.e., $\sup_{Q\in\mathcal{P}(P;\epsilon_{0})}D(Q,P)=\epsilon_{0}\tau_{\infty}^{1/2}$.
This result suggests that, when at most a proportion $\epsilon_{0}$ of data is corrupted, setting $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ will ensure quantitative robustness to $\mathcal{P}(P;\epsilon_{0})$. Note that using a similar argument as in the proof of Proposition 5, Huber’s contamination models can also be related to the MMD balls; see Section B.2.
One practical issue is that $\tau_{\infty}$, the supremum of the Stein kernel, does not always admit a closed-form expression. One way to compute it is by numerical optimization of $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$, but this assumes the optimizer converges and requires extra computation, thus not suitable in high dimension or when evaluation of the score function is costly. We propose an alternative approach, where $\tau_{\infty}$ is approximated by the maximum of the Stein kernel evaluated at the observed data, i.e., $\max_{i=1,\ldots,n}u_{p}(\mathbf{X}_{i},\mathbf{X}_{i})$. This approach requires no extra computation and gives a reasonable estimate with moderate or large $n$. In Section 5, we use this approach and show empirically that the resulting test is still well-calibrated despite this approximation. Further discussions on how this approximation affects the test performance can be found in Section B.3. Overall, we find that this approach still controls the Type-I error rate, even with a small sample size.
@@
Proposition 5 holds for Huber’s contamination models of the form $Q=(1-\epsilon)P+\epsilon R$ with $\epsilon\leq\epsilon_{0}$. In particular, it holds for any contamination distribution $R$. This agnosticism can be useful in practice, because the exact form of contamination is often unknown. On the other hand, when prior knowledge about $R$ is available, it is also possible to incorporate it into the proposed test. For example, if $R$ is known to have a support bounded by some $B>0$, then the bound in Proposition 5 can be tightened by replacing the worst-case bound $\tau_{\infty}=\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})$ with the localized version $\tau_{\infty}=\sup_{\|\mathbf{x}\|_{2}\leq B}u_{p}(\mathbf{x},\mathbf{x})$. As discussed in Fauß et al., (2021), such assumptions are realistic in, e.g., highly regulated experimental environments, where any outliers are known to lie within bounded regions.
Another uncertainty model commonly studied in the literature is the density-band model (Kassam,, 1981; Hafner,, 1993), defined as distributions whose density function lies within an error band of a nominal model with density $p$, i.e., $\{Q\in\mathcal{P}(\mathbb{R}^{d}):\;Q\textrm{ has density }q\textrm{ with }|q(\mathbf{x})-p(\mathbf{x})|\leq\delta(\mathbf{x})\textrm{ for all }\mathbf{x}\}$, for some function $\delta:\mathbb{R}^{d}\to[0,\infty)$. It can be shown that density-band models can be rewritten as Huber’s contamination models (6) with additional constraints on the outlier distribution (Fauß and Zoubir,, 2016). However, compared with Huber’s models, density-band models have the advantage of being more interpretable and more natural for certain forms of model disparity such as heavy tails. The following result suggests how the uncertainty radius $\theta$ should be chosen for such models and is proved in Section A.6.3.
@@
Suppose $k$ satisfies the conditions in Lemma 2 and $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$. Furthermore, assume $Q,P\in\mathcal{P}(\mathbb{R}^{d})$ admit positive densities $q,p$ on $\mathbb{R}^{d}$ and $p\in\mathcal{C}^{1}$. If $|q(\mathbf{x})-p(\mathbf{x})|\leq\delta(\mathbf{x})$ for some function $\delta:\mathbb{R}^{d}\to[0,\infty)$ such that $\delta_{0}\coloneqq\int_{\mathbb{R}^{d}}\delta(\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}<\infty$, then $D(Q,P)\leq\delta_{0}\tau_{\infty}^{1/2}$.
The bound in Proposition 6 is not tight, as its proof involves bounding an integrand that can take negative values by its absolute value. Nevertheless, as we will show in Section 5.1.4, this bound is not overly loose and can still be useful. In particular, it allows us to design a KSD test that is robust to tail misspecification while still maintaining non-trivial power.
Proposition 5 and 6 also reveal a limitation of our robust KSD test—it can be overly conservative to alternatives that are not the intended contamination but still lie within the chosen KSD-ball uncertainty set. For example, to ensure robustness to Huber’s contamination models (6) with tolerance $\epsilon_{0}$, Proposition 5 suggests setting the uncertainty radius to $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$. However, the resulting KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ contains not only Huber’s models, but also other distributions such as density-band models. As a result, the robust KSD test will control Type-I errors uniformly over this ball, regardless of whether they are the intended contamination type. This limitation is not unique to our method, but a generic drawback of all robust tests based on uncertainty sets (Fauß et al.,, 2021). Nevertheless, we will demonstrate in our numerical experiments that, for alternatives outside of the prescribed KSD-ball, our robust test does achieve non-trivial test power.
@@ 5 Numerical Experiments
[…]
@@ 5.1 Toy Gaussian Model
We first consider a Gaussian model $P=\mathcal{N}(0,1)$, in which case the score function of the model is $\mathbf{s}_{p}(x)=-x$ and is hence unbounded. This toy example is simplistic and not representative of the unnormalized models our test is most suited for, but it will nonetheless be helpful to study our algorithmic choices, to verify Theorem 1 and Theorem 3 numerically, and to compare against alternative robust tests.
@@ 5.1.1 Decay Rate of Weighting Function
We first draw random samples $\mathbb{X}_{n}$ of size $n=500$ from $Q=(1-\epsilon)P+\epsilon\delta_{z}$, for different values of $\epsilon\in[0,1]$ and $z\in\mathbb{R}$. This setting mimics the presence of outliers at $z$. Given a stationary kernel $h$, a natural question is “how does the choice of weight $w$ affect the Stein kernel and the test power?”. On the left-hand side of Figure 2, we plot the Stein kernel $x\mapsto u_{p}(x,x)$ for different values of $b$. As $b$ grows, the tails of the Stein kernel are progressively down-weighted. We then plot the rejection probability of the standard KSD tests using these Stein kernels when the outlier is $z=10$ on the right-hand side of Figure 2. As $b$ increases, the rejection probability decreases, suggesting that the test power is lower when the tails of the Stein kernel are overly down-weighted. This is not surprising since the Stein kernel decays faster for larger values of $b$, making the test insensitive to model deviations at the tails. Ideally, $w$ should decay just enough so that $u_{p}$ remains bounded, but not too fast as it would lose power. This highlights the trade-off between robustness and power in the choice of $w$. A similar trade-off has been observed in robust estimation, where an overly decaying $w$ can impede the estimation accuracy (Barp et al.,, 2019, Appendix F.2). Henceforth, we choose $b=1/2$ to balance this trade-off, as it is the smallest value that ensures $u_{p}$ is bounded for sub-Gaussian models.
@@ 5.1.2 Kernel Bandwidth
We run the robust-KSD test with various choices of bandwidths $\lambda^{2}\in\Lambda\cup\{\lambda_{\mathrm{med}}^{2}\}$, where $\Lambda=\{0.01,0.1,1,10,100\}$ and the uncertainty radius $\theta$ is set to control at most $\epsilon=0.05$ contaminations in the sample. We also compare it with the KSDAgg test of Schrab et al., (2022) using the same tilted kernel and aggregating over the collection of bandwidths $\Lambda$. Other configurations of KSDAgg follow the recommendations in Schrab et al., (2022, Section 4.2). This experiment is repeated 200 times to reduce randomness, and the results are reported in Figure 2. For any choice of $\lambda$, the robust-KSD test is able to control the Type-I error when $\epsilon\leq\epsilon_{0}$, where $\epsilon$ is the proportion of contamination in the data, suggesting that robust-KSD remains well-calibrated for all bandwidths. If instead $\epsilon>\epsilon_{0}$, all robust-KSD tests eventually reject $H^{\mathrm{C}}_{0}$, with $\lambda$ chosen by median heuristic achieving the highest power. The KSDAgg test is not robust and more sensitive to contamination than all robust-KSD tests at all contamination levels $\epsilon$. This is not surprising given that KSDAgg is designed to maximize the test power over multiple bandwidths, rendering this test more sensitive to a wide range of alternatives, including contaminated models. As a result, we henceforth do not include KSDAgg in our experiments. A similar study for the standard KSD test is given in Section B.4.
@@ 5.1.3 Standard versus Robust KSD Tests
Using the same data set, we then compare KSD tests with a stationary kernel to tilted-KSD tests and our proposed robust-KSD test in Figure 3. As shown in the left two plots, with a contamination ratio $\epsilon$ of $0.01$ or $0.05$, the standard IMQ-KSD test rejects the point null with high probability for large outlier values $z$. This aligns with the lack of qualitative robustness of tests based on stationary kernels proved in Theorem 1. In contrast, the standard Tilted-KSD test rejects with lower probability for all $z$ values, aligning with the qualitative robustness result with tilted kernels in Theorem 3. Notably, our theoretical results in both Section 3 and Section 4 apply only to fixed kernels, thus excluding kernels based on observed data such as those selected via median heuristic. However, our numerical results suggest that the conclusion of our theory may hold more broadly.
@@
The Type-I error of the tilted-KSD test still exceeds the nominal level $\alpha=0.05$, suggesting that a tilted kernel alone is not enough to enforce quantitative robustness. Notably, the rejection probability of Tilted-KSD is highest when $z=1$ but declines for larger $z$. This is because $u_{p}(z,z)$ with the tilted kernel peaks at around $|z|\approx 1$ and then decreases with $z$ (see Figure 2), making the test more sensitive when $|z|\approx 1$ but less sensitive to large outliers. In contrast, running our robust-KSD test with a pre-set contamination control $\epsilon_{0}=0.05$, we can see from the left two plots of Figure 3 that it remains well-calibrated for all values of outlier $z$. The results also suggest that R-KSD is conservative since its rejection probability is close to 0; we attribute this to the fact that the robust tests are designed to control all types of contamination, of which the injected outlier is only one specific type. To demonstrate the non-trivial power of the robust tests, we show in the right two plots the rejection probability against different contamination ratios $\epsilon$. As $\epsilon$ grows, R-KSD eventually is capable of rejecting the null hypothesis $\mathcal{H}_{0}:Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ with test power approaching one.
@@
We also compared our tests with two existing robust kernel tests, namely the robust MMD tests of Schrab and Kim, (2024) (denoted dcMMD) and of Sun and Zou, (2023, Eq. (38)) (denoted MMD-Dev). Both tests are provably quantitatively robust to Huber’s contamination models, but they are two-sample tests that require extra samples from the model $P$. When the same number of samples from $Q$ and from $P$ are used and the cost of simulation is negligible, the cost of a single evaluation of MMD comparable to that of KSD. However, this cost could be prohibitively expensive when $P$ is a more complex model. Implementation details of these tests are deferred to Section B.2. As shown in Figure 3, both tests are more conservative than our proposed KSD tests. The conservativeness of dcMMD is because it is designed for a contamination model that is stronger than Huber’s model, while the conservativeness of MMD-Dev is because it uses a deviation inequality to construct the decision threshold, which is known to be conservative; see, e.g., Gretton et al., (2012).

### v4

@@ 4.2 Choosing the Uncertainty Radius
Firstly, suppose we are considering Huber’s contamination model $\mathcal{P}(P;\epsilon_{0})$ for some $\epsilon_{0}\in[0,1]$. Then $\theta$ should be chosen such that the KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ contains $\mathcal{P}(P;\epsilon_{0})$. The following result, proved in Section˜A.6, shows how to achieve this given an upper bound on the Stein kernel $\tau_{\infty}=\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})$.
@@
Suppose $k$ satisfies the conditions in Lemma˜2 and $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$. Let $\epsilon_{0}\in[0,1]$. Then $\mathcal{P}(P;\epsilon_{0})\subseteq\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ if $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ and this bound is tight, i.e., $\sup_{Q\in\mathcal{P}(P;\epsilon_{0})}D(Q,P)=\epsilon_{0}\tau_{\infty}^{1/2}$.
This result suggests that, when at most a proportion $\epsilon_{0}$ of data is corrupted, setting $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ will ensure quantitative robustness to $\mathcal{P}(P;\epsilon_{0})$. Note that using a similar argument as in the proof of Proposition˜5, Huber’s contamination models can also be related to the MMD balls; see Section˜B.2.
One practical issue is that $\tau_{\infty}$, the supremum of the Stein kernel, does not always admit a closed-form expression. One way to compute it is by numerical optimization of $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$, but this assumes the optimizer converges and requires extra computation, thus not suitable in high dimension or when evaluation of the score function is costly. We propose an alternative approach, where $\tau_{\infty}$ is approximated by the maximum of the Stein kernel evaluated at the observed data, i.e., $\max_{i=1,\ldots,n}u_{p}(\mathbf{X}_{i},\mathbf{X}_{i})$. This approach requires no extra computation and gives a reasonable estimate with moderate or large $n$. In Section˜5, we use this approach and show empirically that the resulting test is still well-calibrated despite this approximation. Further discussions on how this approximation affects the test performance can be found in Section˜B.3, where we demonstrate that this approach still controls the Type-I error rate, even with a small sample size.
@@
Proposition˜5 holds for Huber’s contamination models of the form $Q=(1-\epsilon)P+\epsilon R$ with $\epsilon\leq\epsilon_{0}$. In particular, it holds for any contamination distribution $R$. This agnosticism can be useful in practice, because the exact form of contamination is often unknown. On the other hand, when prior knowledge about $R$ is available, it is also possible to incorporate it into the proposed test. For example, if $R$ is known to have a support bounded by some $B>0$, then the bound in Proposition˜5 can be tightened by replacing the worst-case bound $\tau_{\infty}=\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})$ with the localized version $\tau_{\infty}=\sup_{\|\mathbf{x}\|_{2}\leq B}u_{p}(\mathbf{x},\mathbf{x})$. As discussed in Fauß et al., (2021), such assumptions are realistic in, e.g., highly regulated experimental environments, where any outliers are known to lie within bounded regions.
[…]
@@
Suppose $k$ satisfies the conditions in Lemma˜2 and $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$. Furthermore, assume $Q,P\in\mathcal{P}(\mathbb{R}^{d})$ admit positive densities $q,p$ on $\mathbb{R}^{d}$ and $p\in\mathcal{C}^{1}$. If $|q(\mathbf{x})-p(\mathbf{x})|\leq\delta(\mathbf{x})$ for some function $\delta:\mathbb{R}^{d}\to[0,\infty)$ such that $\delta_{0}\coloneqq\int_{\mathbb{R}^{d}}\delta(\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}<\infty$, then $D(Q,P)\leq\delta_{0}\tau_{\infty}^{1/2}$.
The bound in Proposition˜6 is not tight, as its proof involves bounding an integrand that can take negative values by its absolute value. Nevertheless, as we will show in Section˜5.1.4, this bound is not overly loose and can still be useful. In particular, it allows us to design a KSD test that is robust to tail misspecification while still maintaining non-trivial power.
Proposition˜5 and 6 also reveal a limitation of our robust KSD test—it can be overly conservative to alternatives that are not the intended contamination but still lie within the chosen KSD-ball uncertainty set. For example, to ensure robustness to Huber’s contamination models (6) with tolerance $\epsilon_{0}$, Proposition˜5 suggests setting the uncertainty radius to $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$. However, the resulting KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ contains not only Huber’s models, but also other distributions such as density-band models. As a result, the robust KSD test will control Type-I errors for all these distributions, regardless of whether they are the intended contamination type. This limitation is not unique to our method; it is a generic drawback of all robust tests based on uncertainty sets (Fauß et al.,, 2021).
@@ 5 Numerical Experiments
We will now evaluate the proposed GOF tests using both synthetic and real data. Unless otherwise mentioned, all standard KSD tests are based on an IMQ kernel $k(\mathbf{x},\mathbf{x}^{\prime})=h_{\mathrm{IMQ}}(\mathbf{x}-\mathbf{x}^{\prime})$ where $h_{\mathrm{IMQ}}(\mathbf{u})=(1+\|\mathbf{u}\|_{2}^{2}/\lambda^{2})^{-1/2}$ with a bandwidth $\lambda^{2}>0$ selected via the median heuristic, i.e., $\lambda_{\mathrm{med}}\;=\;\mathrm{Median}\big{\{}\|\mathbf{X}_{i}-\mathbf{X}_{j}\|_{2}:\;1\leq i<j\leq n\big{\}}$. All tilted-KSD and robust-KSD tests are based on a tilted IMQ kernel with weight $w(\mathbf{x})=(1+\|\mathbf{x}-\mathbf{a}\|_{2}^{2}/c)^{-b}$, where $\mathbf{a}\in\mathbb{R}^{d}$ and $c>0$. Intuitively, $\mathbf{a}$ and $c$ respectively centers and scales the input. We fix $\mathbf{a}=0$ and $c=1$ in all experiments, as all data will always be centered and on a suitable scale. More generally, we could replace $\|\mathbf{x}-\mathbf{a}\|_{2}^{2}/c$ by a weighted norm of the form $(\mathbf{x}-\mathbf{a})^{\top}C(\mathbf{x}-\mathbf{a})$, where $C\in\mathbb{R}^{d\times d}$ is a pre-conditioning matrix, chosen possibly as the empirical covariance matrix or robust estimates of it. Since our experiments will focus on sub-Gaussian models, we choose $b=1/2$. This ensures the Stein kernel is bounded.
@@ 5.1 Toy Gaussian Model
We first consider a Gaussian model $P=\mathcal{N}(0,1)$, in which case the score function of the model is $\mathbf{s}_{p}(x)=-x$ and is hence unbounded. This toy example is simplistic and not representative of the unnormalized models our test is most suited for, but it will nonetheless be helpful to study our algorithmic choices, to verify Theorem˜1 and Theorem˜3 numerically, and to compare against alternative robust tests.
@@ 5.1.1 Decay Rate of Weighting Function
We first draw random samples $\mathbb{X}_{n}$ of size $n=500$ from $Q=(1-\epsilon)P+\epsilon\delta_{z}$, for different values of $\epsilon\in[0,1]$ and $z\in\mathbb{R}$. This setting mimics the presence of outliers at $z$. Given a stationary kernel $h$, a natural question is “how does the choice of weight $w$ affect the Stein kernel and the test power?”. On the left-hand side of Figure˜2, we plot the Stein kernel $x\mapsto u_{p}(x,x)$ for different values of $b$. As $b$ grows, the tails of the Stein kernel are progressively down-weighted. We then plot the rejection probability of the standard KSD tests using these Stein kernels when the outlier is $z=10$ on the right-hand side of Figure˜2. As $b$ increases, the rejection probability decreases, suggesting that the test power is lower when the tails of the Stein kernel are overly down-weighted. This is not surprising since the Stein kernel decays faster for larger values of $b$, making the test insensitive to model deviations at the tails. Ideally, $w$ should decay just enough so that $u_{p}$ remains bounded, but not too fast as it would lose power. This highlights the trade-off between robustness and power in the choice of $w$. A similar trade-off has been observed in robust estimation, where an overly decaying $w$ can impede the estimation accuracy (Barp et al.,, 2019, Appendix F.2). Henceforth, we choose $b=1/2$ to balance this trade-off, as it is the smallest value that ensures $u_{p}$ is bounded for sub-Gaussian models.
@@ 5.1.2 Kernel Bandwidth
We run the robust-KSD test with various choices of bandwidths $\lambda^{2}\in\Lambda\cup\{\lambda_{\mathrm{med}}^{2}\}$, where $\Lambda=\{0.01,0.1,1,10,100\}$ and the uncertainty radius $\theta$ is set to control at most $\epsilon=0.05$ contaminations in the sample. We also compare it with the KSDAgg test of Schrab et al., (2022) using the same tilted kernel and aggregating over the collection of bandwidths $\Lambda$. Other configurations of KSDAgg follow the recommendations in Schrab et al., (2022, Section 4.2). This experiment is repeated 200 times to reduce randomness, and the results are reported in Figure˜2. For any choice of $\lambda$, the robust-KSD test is able to control the Type-I error when $\epsilon\leq\epsilon_{0}$, where $\epsilon$ is the proportion of contamination in the data, suggesting that robust-KSD remains well-calibrated for all bandwidths. If instead $\epsilon>\epsilon_{0}$, all robust-KSD tests eventually reject $H^{\mathrm{C}}_{0}$, with $\lambda$ chosen by median heuristic achieving the highest power. The KSDAgg test is not robust and more sensitive to contamination than all robust-KSD tests at all contamination levels $\epsilon$. This is not surprising given that KSDAgg is designed to maximize the test power over multiple bandwidths, rendering this test more sensitive to a wide range of alternatives, including contaminated models. As a result, we henceforth do not include KSDAgg in our experiments. A similar study for the standard KSD test is given in Section˜B.4.
@@ 5.1.3 Standard versus Robust KSD Tests
Using the same data set, we then compare KSD tests with a stationary kernel to tilted-KSD tests and our proposed robust-KSD test in Figure˜3. As shown in the left two plots, with a contamination ratio $\epsilon$ of $0.01$ or $0.05$, the standard IMQ-KSD test rejects the point null with high probability for large outlier values $z$. This aligns with the lack of qualitative robustness of tests based on stationary kernels proved in Theorem˜1. In contrast, the standard Tilted-KSD test rejects with lower probability for all $z$ values, aligning with the qualitative robustness result with tilted kernels in Theorem˜3. Notably, our theoretical results in both Section˜3 and Section˜4 apply only to fixed kernels, thus excluding kernels based on observed data such as those selected via median heuristic. However, our numerical results suggest that the conclusion of our theory may hold more broadly.
@@
The Type-I error of the tilted-KSD test still exceeds the nominal level $\alpha=0.05$, suggesting that a tilted kernel alone is not enough to enforce quantitative robustness. Notably, the rejection probability of Tilted-KSD is highest when $z=1$ but declines for larger $z$. This is because $u_{p}(z,z)$ with the tilted kernel peaks at around $|z|\approx 1$ and then decreases with $z$ (see Figure˜2), making the test more sensitive when $|z|\approx 1$ but less sensitive to large outliers. In contrast, running our robust-KSD test with a pre-set contamination control $\epsilon_{0}=0.05$, we can see from the left two plots of Figure˜3 that it remains well-calibrated for all values of outlier $z$. The results also suggest that R-KSD is conservative since its rejection probability is close to 0; we attribute this to the fact that the robust tests are designed to control all types of contamination, of which the injected outlier is only one specific type. To demonstrate the non-trivial power of the robust tests, we show in the right two plots the rejection probability against different contamination ratios $\epsilon$. As $\epsilon$ grows, R-KSD eventually is capable of rejecting the null hypothesis $\mathcal{H}_{0}:Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ with test power approaching one.
@@
We also compared our tests with two existing robust kernel tests, namely the robust MMD tests of Schrab and Kim, (2024) (denoted dcMMD) and of Sun and Zou, (2023, Eq. (38)) (denoted MMD-Dev). Both tests are provably quantitatively robust to Huber’s contamination models, but they are two-sample tests that require extra samples from the model $P$. When the same number of samples from $Q$ and from $P$ are used and the cost of simulation is negligible, the cost of a single evaluation of MMD comparable to that of KSD. However, this cost could be prohibitively expensive when $P$ is a more complex model. Implementation details of these tests are deferred to Section˜B.2. As shown in Figure˜3, both tests are more conservative than our proposed KSD tests. The conservativeness of dcMMD is because it is designed for a contamination model that is stronger than Huber’s model, while the conservativeness of MMD-Dev is because it uses a deviation inequality to construct the decision threshold, which is known to be conservative; see, e.g., Gretton et al., (2012).